The length of the diagonal of a television screen is the measurement used to report its size. Please Calculate the size of a television, to the nearest centimetre, where the screen measures 56 cm by 40 cm.
Answer by fractalier(6550)

Given that the screen is a rectangle, the diagonal is just the hypotenuse of a right triangle whose legs are 56 and 40...
so we use the Pythagorean Theorem...
a^2 + b^2 = c^2
56^2 + 40^2 = c^2
3136 + 1600 = c^2
4736 = c^2 so that
c is about 68.8 or 69 to the nearest cm